solution of the given equations Find the general Y - 3y = - 72 5ın 6oc - 72 cos loc + 3b es
Answer #1

62 solution: Given y-364 = -72 sin 6x - 72 Cos 6x +36e Let y = A (70) be a trial solution equation becomes Then the auxilia: y = 3x0+ sinore & Cosor Hence the general solution of the given differential equation is 64 y = qe tege 4 ne 4 sinte tlosbelike please.